OpenStudy (anonymous):

Use the method of elimination to solve the system 3x+3y=-9 -3x+1y=-3 Your answer is (x,y)= ( , ) I don't understand this question? 😭

OpenStudy (anonymous):

Do you understand anything about this question?

OpenStudy (mathmale):

Here, "elimination" means to get rid of (eliminate) one variable at a time. In this case the given problem is very straightforward. Copy down this system of equations: 3x+3y=-9 -3x+1y=-3 and then add the terms together vertically. In other words, add: 3x -3x ______

OpenStudy (anonymous):

No not at all

OpenStudy (mathmale):

Then what? Solve this simplified equation for y. Then find x. Please follow these instructions, then ask for specific help if you need it.

OpenStudy (anonymous):

Y = -3 I got y but I still can't get x

OpenStudy (mathmale):

Substitute y=-3 into either of the given equations. This will leave you with only one variable: x Solve for x. Show your work, please.